OpenStudy (anonymous):

The sum of the first and third consecutive integers is equal to twice the second. What are the Integers? Explain.

OpenStudy (anonymous):

let the consecutive integers be x , x+1 , x+2 According to the question : \[x+(x+2) = 2(x+1)\] \[x+x+2=2x+2\] \[2x+2=2x+2\] i do think it is a logic ,,,,,,,,,,,,,,,,, all the consecutive integers accept in this case i mean : 5,6,7 : 5+7=2(6) 12=12 6,7,8 : 6+8 = 2(7) => 14=14

OpenStudy (anonymous):

\(x+x+2=2(x+1) \implies 2(x+1)=2(x+1) \implies x=x\). That means it's applicable to any three consecutive numbers. Take for example 2,3,4 or 5,6,7.

OpenStudy (anonymous):

You can see that \(2+4=2\times3\) and that \(5+7=2\times6\) and so on.

OpenStudy (anonymous):

-5,-4,-3 -5+(-3) = 2(-4) -5-3=-8 -8=-8 anwar 's answer agrees with mine too so now i am confirmed that all the integers are applicable in these situations ( consecutive )

OpenStudy (anonymous):

WHETHER THEY ARE POSITIVE OR NEGATIVE

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

\[\text{nppbleh789}\]